OpenStudy (richie63):

Easy! And Quick! I will give a medal and best Response. Which action will solve the equation? m – 20 = 9 (Points : 1) Add 9 to each side. Subtract 9 from each side. Add 20 to each side. Subtract 20 from each side

OpenStudy (anonymous):

add 20

OpenStudy (anonymous):

Correct Add 20 to each side. Since it's got the equal sign in the middle what you do to the left side you need to do to the right side. And you need to do the inverse to get M alone. Hope I helped :D

OpenStudy (anonymous):

Yes, @xxbabi3gurlxx is right. It is SUBTRACT 20 FROM BOTH SIDES.

OpenStudy (anonymous):

you mean ADD

OpenStudy (richie63):

Solve for x. |x| – 7 = 8 (Points : 1) x = –1 or 1 x = 1 x = –15 or 15 x = 15

OpenStudy (anonymous):

ok add 7 to both sides x-7=8 +7 +7 x=8+7 x=15

OpenStudy (richie63):

Lin parachutes from a plane at a height of 22,000 ft above sea level. Several minutes later she notes from an instrument on her wrist that she had dropped 10,000 ft. How far above sea level is she now? (Points : 1) –12,000 ft 6,000 ft 10,000 ft 12,000 ft Thanks @xxbabi3gurlxx

OpenStudy (anonymous):

Well, if she was at 22,000 ft above sea level and she drops 10,000 feet our equation is 22,000-10,000=? Solve this for the answer.

OpenStudy (richie63):

Thanks! @ohlookidontunderstandanotherthing Which equation models the problem below if c represents the number of kittens whose eyes are closed? There are 22 kittens in a photograph, and 16 of them have their eyes open. How many kittens have their eyes closed? (Points : 1) 16 + c = 22 22 + c = 16 c – 16 = 22 16 – c = 22

OpenStudy (anonymous):

C! Need an explanation or just checking your answers?

OpenStudy (richie63):

@ohlookidontunderstandanotherthing Thanks! ^.^ Your Awesome!

OpenStudy (anonymous):

I try! Thanks! Good luck!!

OpenStudy (richie63):

1. Solve. x – 7 = –19 (Points : 1) -26 -12 12 26

OpenStudy (anonymous):

x=-12

OpenStudy (richie63):

@ohlookidontunderstandanotherthing @xxbabi3gurlxx Thanks! For the Help! You Two Rock!!! ^.^